# Tracing the impacts of Mount Pinatubo eruption on global climate using spatially-varying changepoint detection

### v1

@@ 6 Discussion
Motivated by need to quantify of the impact of the Mt. Pinatubo eruption on global climate and specify its spatial pattern, we proposed a Bayesian framework to simultaneously detect and estimate changepoint for spatio-temporal data. Our approach can identify which spatial locations have a changepoint and, at the same time, allows the changepoints and the magnitude of change to vary spatially, as opposed to assuming that all locations experience the same amount of change at the same time point. Furthermore, our model respects the diffusion pattern of the impact from an SAI event and takes advantage of spatial correlation in changepoint detection and estimation. In particular, our method ensures that changepoints at locations other than the event origin occur strictly after the initial changepoint. This feature helps to prevent detecting changepoints caused by unrelated events, providing a more focused analysis of our desired atmospheric event. The validity and effectiveness of our approach are demonstrated through simulations. By applying our method to AOD and surface temperature data, we successfully captured the spatial patterns of changepoints, which reveals the progression of injected aerosol and its spatially heterogeneous impact on climate.
@@
For convenience, we have opted for the exponential covariance function, a stationary model, to address spatial correlation in the data. However, given our large spatial domain, it may be more realistic to consider a nonstationary covariance function, as suggested by strategies in Shand and Li (2017). We anticipate that computations would be more expensive when considering nonstationary spatial correlation. Additionally, we by no means imply a causal relationship between the Mt. Pinatubo eruption and the detected changepoints. These changepoints could be due to other climate and weather events, such as the El Niño effect, or other internal variations. Our aim in this article is to capture a plausible impact from the Mt. Pinatubo volcanic eruption. Separating Mt. Pinatubo’s impact from other scenarios and establishing a pure causal relationship between the volcanic eruption and climate impact is nontrivial and is a topic we reserve for future studies.
@@
Funding Statement: Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525. SAND2024-XXXX.
7 Appendix
7.1 MCMC sampling
@@
$\boldsymbol{\mu_{\tau}}$ $=\boldsymbol{\mu}_{F}+\boldsymbol{\mu}_{R}\circ\mathbf{1_{\tau}^{+}}$ (11)
$=\mathbf{Z}_{\tau}\boldsymbol{\alpha}+\boldsymbol{\mu}_{R}\circ\mathbf{1_{\tau}^{+}}$ (12)
@@ 1. Step 1 :
      $\widetilde{\pi}_{k}$ $\propto\exp\left(-\frac{1}{2}(\mathbf{Z}_{\tau}\boldsymbol{\alpha}+(\boldsymbol{\epsilon_{\gamma}}+\boldsymbol{\mu}_{R})\circ\mathbf{1_{\tau}^{+}})^{T}(\sigma^{2}_{1}\mathbf{I}_{MN}+\boldsymbol{\Sigma_{U}})^{-1}(\mathbf{Z}_{\tau}\boldsymbol{\alpha}+(\boldsymbol{\epsilon_{\gamma}}+\boldsymbol{\mu}_{R})\circ\mathbf{1_{\tau}^{+}})\right).$
@@
      See section 7.1.1 on how to sample $\boldsymbol{\Delta}$ and $\boldsymbol{\mu}_{R}$
@@ 3. Step 3 :
      $\sigma^{2}_{1}\mid\cdot$ $\sim IG\left(a_{1}+\frac{MN}{2},b_{1}+\frac{(\mathbf{Y}-\mathbf{U}-\boldsymbol{\mu_{\tau}}-\boldsymbol{\epsilon_{\gamma}}\circ\mathbf{1_{\tau}^{+}}))^{T}(\mathbf{Y}-\mathbf{U}-\boldsymbol{\mu_{\tau}}-\boldsymbol{\epsilon_{\gamma}}\circ\mathbf{1_{\tau}^{+}}))}{2}\right)$
@@
      $f(\psi_{\Delta}\mid\cdot)$ $\propto|\mathbf{R}(\psi_{\Delta})|^{-1/2}\exp\left(-\frac{(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})^{T}\mathbf{R}(\psi_{\Delta})^{-1}(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})}{2\sigma^{2}_{\Delta}}\right)\mathds{1}(l_{3}<\psi_{\Delta}<u_{3})$
@@
      $r_{\psi_{\Delta}}$ $=\left(\frac{|\mathbf{R}(\psi_{\Delta}^{*})|}{|\mathbf{R}(\psi_{\Delta}^{k-1}|}\right)^{-1/2}\exp\left(-\frac{(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})^{T}(\mathbf{R}(\psi_{\Delta}^{*})^{-1}-\mathbf{R}(\psi_{\Delta}^{k-1})^{-1})(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})}{2\sigma^{2}_{\Delta}}\right)$
@@
7.1.1 Adaptive Metropolis Hastings
@@
We sample $\boldsymbol{\Delta}=(\Delta_{s_{1}},...,\Delta_{s_{N}})^{T}$ using adaptive component-wise Metropolis Hastings algorithm Roberts and Rosenthal (2009) as follows:
@@
        Draw $\log\Delta^{*}_{s_{i}}$ from proposal density $q(\log\Delta^{*}_{s_{i}}\mid\log\Delta^{k-1}_{s_{i}})\sim N(\log\Delta_{s_{i}}^{k-1},\exp(ls_{i}))$
@@
        $(\displaystyle r=\exp(-$$(\log\boldsymbol{\Delta}^{*}-\log\boldsymbol{\Delta}^{k-1})^{T}\mathbf{R}(\psi_{\Delta})^{-1}(\log\boldsymbol{\Delta}^{*}+\log\boldsymbol{\Delta}^{k-1}-\mathbf{X}\boldsymbol{\beta})/2\sigma^{2}_{\Delta}+$
        $(\boldsymbol{\mu_{\tau}}^{*}+\boldsymbol{\mu_{\tau}}^{k-1}+\boldsymbol{\epsilon_{\gamma}}\circ(\mathbf{1_{\tau}^{+}}^{*}+\mathbf{1_{\tau}^{+}}^{k-1})-2(\mathbf{Y}-\mathbf{U}))^{T}(\boldsymbol{\mu_{\tau}}^{*}-\boldsymbol{\mu_{\tau}}^{k-1}+\boldsymbol{\epsilon_{\gamma}}\circ(\mathbf{1_{\tau}^{+}}^{*}-\mathbf{1_{\tau}^{+}}^{k-1})/2\sigma^{2}_{1})),$
@@
[…]
@@
$(\displaystyle r=\exp(-$$(\boldsymbol{\gamma_{0R}}^{*}+\boldsymbol{\gamma_{0R}}^{k-1})^{T}\boldsymbol{\Sigma_{\gamma_{0}}}^{-1}(\boldsymbol{\gamma_{0R}}^{*}-\boldsymbol{\gamma_{0R}}^{k-1})/2+$
@@
$(\displaystyle r=.\exp(-$$(\boldsymbol{\gamma_{1R}}^{*}+\boldsymbol{\gamma_{1R}}^{k-1})^{T}\boldsymbol{\Sigma_{\gamma_{1}}}^{-1}(\boldsymbol{\gamma_{1R}}^{*}-\boldsymbol{\gamma_{1R}}^{k-1})/2+$
@@
7.2 Table
@@
$56.0^{\circ}$ Jul 1991 0 (Jun 1991, Aug 1993)
$48.5^{\circ}$ Jul 1991 0 (Jun 1991, Mar 1992)
$41.0^{\circ}$ Aug 1991 0 (Jul 1991, Feb 1992)
$33.5^{\circ}$ Sep 1991 0 (Jul 1991, Nov 1991)
$26.0^{\circ}$ Oct 1991 0.020 (Jun 1991, Aug 1994)
$18.5^{\circ}$ Dec 1996 (No CP) 0.242 (Sep 1991, Dec 1996)
$11.0^{\circ}$ Dec 1996 (No CP) 0.979 (Dec 1995, Dec 1996)
$3.5^{\circ}$ Dec 1996 (No CP) 0.960 (Jan 1996, Dec 1996)
$4.0^{\circ}$ Dec 1996 (No CP) 0.958 (Dec 1995, Dec 1996)
$11.5^{\circ}$ Sep 1991 0 (Jul 1991, Oct 1991)
$19.0^{\circ}$ Sep 1991 0 (Jun 1991, Nov 1991)
$26.5^{\circ}$ Dec 1996 (No CP) 0.976 (Dec 1996, Dec 1996)
$34.0^{\circ}$ Dec 1996 (No CP) 0.987 (Dec 1996, Dec 1996)
$41.5^{\circ}$ Jun 1992 0 (Sep 1991, Oct 1992)
$49.0^{\circ}$ Dec 1996 (No CP) 0.817 (Mar 1992, Dec 1996)
$56.5^{\circ}$ Dec 1996 (No CP) 0.818 (Jan 1992, Dec 1996)
Table 3: Summary of detected changepoints

### v2

@@ 6 Discussion
Motivated by the need to quantify the impact of the Mt. Pinatubo eruption on regional climate and specify its spatial pattern, we proposed a Bayesian framework to simultaneously detect and estimate changepoint for spatio-temporal data. Our approach can identify which spatial locations have a changepoint and, meantime, allows the changepoints and the magnitude of change to vary spatially, as opposed to assuming that all locations experience the same amount of change at the same time point. Furthermore, our model respects the diffusion pattern of the impact of the volcanic eruption and takes advantage of spatial correlation in changepoint detection and estimation. In particular, our method ensures that changepoints at locations other than the event origin occur strictly after the initial changepoint. This feature helps to prevent detecting changepoints caused by unrelated events, providing a more focused analysis of the targeted atmospheric event. The validity and effectiveness of our approach are demonstrated through simulations. By applying our method to AOD and surface temperature data, we successfully captured the spatial patterns of changepoints, which reveals the progression of injected aerosol and its spatially heterogeneous impact on regional aerosol optical depth and surface temperatures.
For convenience, we have opted for the exponential covariance function, a stationary model, to address spatial correlation in the data. However, given our large spatial domain, it may be more realistic to consider a nonstationary covariance function, such as those proposed by Shand and Li (2017). Other possible improvements of covariance models include employing a chordal of circular Matérn covariance function instead of an exponential covariance model (Guinness and Fuentes, 2016), and a more flexible nonseparable or even asymmetric space-time covariance model (Gneiting, 2002) for $U({\bf s},t)$. Another interesting extension of our work is to directly consider a bivariate time series of AOD and surface temperature in changepoint detection and estimation. This may especially benefit the surface temperature due to the weaker signal in temperature than in AOD. This extension is expected to be involving, though, as the dependency structure of the bivariate data will be complex, and the changepoint processes for the two variables will also interact with each other. Additionally, our current method focuses on detecting the time at which the impact of the eruption first reaches each location. Capturing the whole life process of Pinatubo impact including when the data “returns to normal” will be an exciting but nontrial extention of our current method as that requires estimation of more than one changepoint at each location, aligning with the concept of “epidemic changepoints” (Tucker and Yarger, 2023).
@@
Finally, we by no means imply a causal relationship between the Mt. Pinatubo eruption and the detected changepoints. These changepoints could be due to other climate and weather events, such as the El Niño effect, or other internal variations. Our aim in this article is to capture a plausible impact of the Mt. Pinatubo volcanic eruption. Separating Mt. Pinatubo’s impact from other scenarios and establishing a pure causal relationship between the volcanic eruption and climate impact is nontrivial and is a topic we reserve for future studies.
Acknowledgments
Funding Statement
@@ Acknowledgments
This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525. SAND2024-XXXX. Li’s research is also partially supported by NSF-2124576.
Appendix A Supplement to “Tracing the impacts of Mount Pinatubo eruption on regional climate using spatially-varying changepoint detection”
A.1 MCMC sampling
@@
$\boldsymbol{\mu}_{\tau}$ $=\boldsymbol{\mu}_{F}+\boldsymbol{\mu}_{R}\circ\mathbf{1_{\tau}^{+}}$
$=\mathbf{Z}_{\tau}\boldsymbol{\alpha}+\boldsymbol{\mu}_{R}\circ\mathbf{1_{\tau}^{+}}$
@@ 1. Step 1 :
      $\log\widetilde{\pi}_{k}\propto-\frac{1}{2}$ $(\mathbf{Z}_{\tau}\boldsymbol{\alpha}+(\boldsymbol{\epsilon}_{\gamma}+\boldsymbol{\mu}_{R})\circ\mathbf{1_{\tau}^{+}})^{T}(\sigma^{2}_{1}\mathbf{I}_{MN}+\boldsymbol{\Sigma_{U}})^{-1}(\mathbf{Z}_{\tau}\boldsymbol{\alpha}+(\boldsymbol{\epsilon}_{\gamma}+\boldsymbol{\mu}_{R})\circ\mathbf{1_{\tau}^{+}}).$
@@
      See section A.1.1 on how to sample $\boldsymbol{\Delta}$ and $\boldsymbol{\mu}_{R}$
@@ 3. Step 3 :
      $\sigma^{2}_{1}\mid\cdot$ $\sim IG\left(a_{1}+\frac{MN}{2},b_{1}+\frac{(\mathbf{Y}-\mathbf{U}-\boldsymbol{\mu}_{\tau}-\boldsymbol{\epsilon}_{\gamma}\circ\mathbf{1_{\tau}^{+}})^{T}(\mathbf{Y}-\mathbf{U}-\boldsymbol{\mu}_{\tau}-\boldsymbol{\epsilon}_{\gamma}\circ\mathbf{1_{\tau}^{+}})}{2}\right)$
@@
      $f(\psi_{\Delta}\mid\cdot)$ $\propto|\mathbf{R}(\psi_{\Delta})|^{-1/2}\exp\left(-\frac{(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})^{T}\mathbf{R}(\psi_{\Delta})^{-1}(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})}{2\sigma^{2}_{\Delta}}\right)$
      $\cdot\mathds{1}(l_{3}<\psi_{\Delta}<u_{3})$
@@
      $r_{\psi_{\Delta}}$ $=\left(\frac{|\mathbf{R}(\psi_{\Delta}^{*})|}{|\mathbf{R}(\psi_{\Delta}^{k-1}|}\right)^{-\frac{1}{2}}$
      $\phantom{=}\cdot\exp\left(-\frac{(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})^{T}(\mathbf{R}(\psi_{\Delta}^{*})^{-1}-\mathbf{R}(\psi_{\Delta}^{k-1})^{-1})(\log\boldsymbol{\Delta}-\mathbf{X}\boldsymbol{\beta})}{2\sigma^{2}_{\Delta}}\right)$
@@
A.1.1 Adaptive Metropolis Hastings
@@
We sample $\boldsymbol{\Delta}=(\Delta_{s_{1}},...,\Delta_{s_{N}})^{T}$ using adaptive component-wise Metropolis Hastings algorithm as follows:
@@
        Draw $\log\Delta^{*}_{s_{i}}$ from proposal density
        $q(\log\Delta^{*}_{s_{i}}\mid\log\Delta^{k-1}_{s_{i}})\sim N(\log\Delta_{s_{i}}^{k-1},\exp(ls_{i})).$
@@
        $r=\exp($- $(\log\boldsymbol{\Delta}^{*}-\log\boldsymbol{\Delta}^{k-1})^{T}\mathbf{R}(\psi_{\Delta})^{-1}(\log\boldsymbol{\Delta}^{*}+\log\boldsymbol{\Delta}^{k-1}-\mathbf{X}\boldsymbol{\beta})/2\sigma^{2}_{\Delta}+$
        $(\boldsymbol{\mu_{\tau}}^{*}+\boldsymbol{\mu_{\tau}}^{k-1}+\boldsymbol{\epsilon}_{\gamma}\circ(\mathbf{1_{\tau}^{+}}^{*}+\mathbf{1_{\tau}^{+}}^{k-1})-2(\mathbf{Y}-\mathbf{U}))^{T}$
        $(\boldsymbol{\mu_{\tau}}^{*}-\boldsymbol{\mu_{\tau}}^{k-1}+\boldsymbol{\epsilon}_{\gamma}\circ(\mathbf{1_{\tau}^{+}}^{*}-\mathbf{1_{\tau}^{+}}^{k-1})/2\sigma^{2}_{1})),$
@@
        Calculate the average acceptance rate for the $j^{\textsuperscript{th}}$ batch as $a_{j}=\sum\limits_{l=k-B+1}^{k}a^{l}/B.$
@@
$r=\exp($ $-(\boldsymbol{\gamma_{0R}}^{*}+\boldsymbol{\gamma_{0R}}^{k-1})^{T}\boldsymbol{\Sigma_{\gamma_{0}}}^{-1}(\boldsymbol{\gamma_{0R}}^{*}-\boldsymbol{\gamma_{0R}}^{k-1})/2+$
@@
$r=.\exp($- $(\boldsymbol{\gamma_{1R}}^{*}+\boldsymbol{\gamma_{1R}}^{k-1})^{T}\boldsymbol{\Sigma_{\gamma_{1}}}^{-1}(\boldsymbol{\gamma_{1R}}^{*}-\boldsymbol{\gamma_{1R}}^{k-1})/2+$
@@
Appendix B Tables and Figures
Fig. 10: Boxplots of RMSE, FPR, FNR, the empirical coverage probability and the length of 95% credible intervals under $\gamma_{0}\in\{1.5,3\}$ and $\sigma^{2}_{\gamma}\in\{3,5\}$ . “ST” is our spatio-temporal model and “1D” is the univariate method.
@@ Appendix B Tables and Figures
$86.0^{\circ}$S Dec 1995 (No CP) 0.9905 (Dec 1995, Dec 1995)
$78.5^{\circ}$S Dec 1995 (No CP) 0.9888 (Dec 1995, Dec 1995)
$71.0^{\circ}$S Dec 1995 (No CP) 0.9762 (Dec 1995, Dec 1995)
$63.5^{\circ}$S Dec 1995 (No CP) 0.7955 (Apr 1992, Dec 1995)
$56.0^{\circ}$S Mar 1992 0.0037 (Jul 1991, Jul 1994)
$48.5^{\circ}$S Oct 1991 0 (Jul 1991, Feb 1992)
$41.0^{\circ}$S Sep 1991 0 (Jul 1991, May 1992)
$33.5^{\circ}$S Sep 1991 0.0001 (Jul 1991, Jan 1992)
$26.0^{\circ}$S Oct 1991 0.0112 (Jun 1991, Nov 1992)
$18.5^{\circ}$S Dec 1995 (No CP) 0.8470 (Apr 1995, Dec 1995)
$11.0^{\circ}$S Sep 1991 0 (Aug 1991, Aug 1992)
$3.5^{\circ}$S Apr 1992 0 (Aug 1991, Sep 1991)
$4.0^{\circ}$N May 1992 0 (Oct 1991, Aug 1992)
$11.5^{\circ}$N Sep 1991 0 (Jul 1991, Dec 1991)
$19.0^{\circ}$N Sep 1991 0 (Jun 1991, Nov 1991)
$26.5^{\circ}$N Dec 1995 (No CP) 0.9878 (Dec 1995, Dec 1995)
$34.0^{\circ}$N Dec 1995 (No CP) 0.9763 (Dec 1995, Dec 1995)
$41.5^{\circ}$N Sep 1991 0 (Jul 1991, Jun 1992)
$49.0^{\circ}$N Sep 1991 0 (Jul 1991, Mar 1992)
$56.5^{\circ}$N Dec 1995 (No CP) 0.9788 (Dec 1995, Dec 1995)
$64.0^{\circ}$N Dec 1995 (No CP) 0.9986 (Dec 1995, Dec 1995)
$71.5^{\circ}$N Dec 1995 (No CP) 0.9984 (Dec 1995, Dec 1995)
$79.0^{\circ}$N Dec 1995 (No CP) 0.9943 (Dec 1995, Dec 1995)
$86.5^{\circ}$N Dec 1995 (No CP) 0.9984 (Dec 1995, Dec 1995)
Table 4: Summary of detected changepoints for surface temperature
Fig. 11: Residual plot of (a) stratospheric AOD and (b) surface temperature under $\sigma^{2}_{1}=\sigma^{2}_{2}$